OpenStudy (anonymous):

Show that the series sin[(-1)^n/n] converges conditionally ?

OpenStudy (anonymous):

i've already proved that lim of it approaches 0, but how can we decide it to be convergent ?

OpenStudy (anonymous):

i mean, convergent then lim = 0 but it's not true vice versa

OpenStudy (anonymous):

if the limit of the sequence approaches 0, then we can not conclude the series is convergent. One example is 1/n. ∑sin( (-1)^n / n) is a alternating series yes?

OpenStudy (anonymous):

yes, it's alternating series, but only the (-1)^n/n is convergent. How can we combine with sine function?

OpenStudy (anonymous):

well, use the fact that sin(-a) = -sin(a) we have ∑sin( (-1)^n / n) =∑ (-1)^n sin(1/n) , yes?

OpenStudy (anonymous):

yes, it approaches 0, but how to prove it's convergent ?

OpenStudy (anonymous):

the criteria to test whether an alternating series is convergent is: 1) a(n+1) <= a(n) 2) lim a(n) = 0 n-> 0

OpenStudy (anonymous):

you already showed the second part. Now you need to show that sin(1/(n+1)) <= sin(1/n)

OpenStudy (anonymous):

so we have 1/(n+1) <= 1/n (which is true) since both criteria meets, by thy Alternating Series Test, the series converge.

OpenStudy (anonymous):

to so the series converge conditionally, show that ∑ |a(n)| does not converge

OpenStudy (anonymous):

okay, i see

OpenStudy (anonymous):

| (-1)^n sin(1/n) | = sin(1/n) compare with (1/n) lim sin(1/n) / (1/n) = 1 n->inf which means both series either converge or diverge. But 1/n diverges, hence, sin(1/n) also diverges

OpenStudy (anonymous):

So the series converge conditionally

OpenStudy (anonymous):

But how can sin(1/n) /(1/n) = 1 ?

OpenStudy (anonymous):

L'hospital rule

OpenStudy (anonymous):

I see, thank you!